OpenStudy (kirbykirby):

Let \(d \in \mathbb{R} \) and let \(X\) be a random variable with support \((0,\infty\)) and \(Y\) be a random variable defined as follows \[ Y = \begin{cases} 0 & , X \le d \\ X & , X > d \end{cases} \] Find the cdf and pdf of Y.

OpenStudy (kirbykirby):

I have the solution written as cdf: \[ F_Y(y) = \begin{cases} F_X(d) & , y=0 \\ F_X(d) &, 0<y<d \\ F_X(y) & , y\ge d \ \end{cases} \] pdf: \[ f_Y(y) = \begin{cases} F_X(d) & , y=0 \\ 0 &, 0<y<d \\ f_X(y) & , y\ge d \ \end{cases} \] I'm hoping for some clarification on how to find these! This is related to mixed random variables (with continuous and discrete parts) which I'm not too familiar with!

OpenStudy (kirbykirby):

I'm guessing it has to do with manipulating the expression \(P(Y\le y)\) though I'm not to clear on it in this case

OpenStudy (zarkon):

\[F_Y(y)=P(Y\le y)=P(Y\le y,X\le d)+P(Y\le y,X>d)\] \[P(Y\le y|X\le d)P(X\le d)+P(Y\le y|X>d)P(X>d)\]